# Mut4All: Fuzzing Compilers via LLM-Synthesized Mutators Learned from Bug Reports

### v1

@@ Abstract.
In this work, we present Mut4All, a fully automated and language-agnostic framework that synthesizes mutation operators by leveraging Large Language Models (LLMs) and compiler-specific knowledge extracted from bug reports on open-source repositories. It is a multi-agent framework comprising three specialized LLM agents: (1) a mutator invention agent, which identifies mutation targets and generates mutator names, signatures, and descriptions based on compiler-related insights (e.g., language features explicitly mentioned in bug reports); (2) a mutator implementation synthesis agent, which is enhanced via fine-tuning to generate initial mutator implementations; and (3) a mutator refinement agent, which verifies and corrects the synthesized mutators using feedback from unit tests. This pipeline analyzes 1000 bug reports (500 each for Rust and C++), and synthesizes 319 and 403 mutators for them. Each mutator only costs approximately \$0.08 using GPT-4o APIs. Using these mutators, our customized fuzzer has reported 62 bugs in the Rust compilers (rustc and gccrs), 38 of which were confirmed as previously unknown, with 7 already fixed by developers. It has also reported 34 bugs in C++ compilers (GCC and Clang), including 16 confirmed as new ones, 1 of which has been resolved. Experimental results also show that Mut4All discovers significantly more unique crashes than existing approaches, and achieves the highest coverage on Rust and the second-highest on C++.
Compiler, Mutation-Based Compiler Fuzzing, Compiler Bugs, Rust, C++
copyright: noneccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Compilers
@@ 1. Introduction
Compiler fuzzing (Chen et al., 2020; Marcozzi et al., 2019; Chen et al., 2016a) has proven effective in improving the reliability of production compilers such as GCC and rustc, uncovering numerous bugs over the past decades. Existing methods can be broadly categorized into two types: generation-based and mutation-based. Generation-based fuzzers, such as Csmith (Yang et al., 2011) and YARPGen (Livinskii et al., 2020, 2023), generate syntactically valid programs from scratch using the compiler’s input grammar. However, ensuring the semantic correctness of these test programs often requires substantial manual effort to design templates, particularly for modern programming languages with intricate features such as ownership and macros, which limits their scalability and adaptability.
In contrast, mutation-based fuzzers such as GrayC (Even-Mendoza et al., 2023) and MetaMut (Ou et al., 2024), which generate new test programs by applying mutation operators (i.e., mutators) to existing real-world programs (i.e., seed programs), are often preferred for compiler fuzzing, as they leverage existing code to avoid the need for designing complex templates and preserve both the syntactic and semantic correctness of test programs. That being said, their effectiveness heavily depends on the quality and expressiveness of the mutators they employ. Moreover, existing mutation-based fuzzers still face several major challenges in mutator design and implementation:
C1: Limited Expressiveness. Most existing mutation-based fuzzers define mutators as compositions of simple operators, such as arithmetic adjustments (e.g., replacing + with -) or control-flow rewrites (e.g., converting switch statements into if-else chains), as exemplified by GrayC (Even-Mendoza et al., 2023) and, to a more advanced extent, MetaMut (Ou et al., 2024). While effective at exposing bugs in basic language constructs, these mutators inherently lack expressiveness—the ability to construct program variants that incorporate advanced language features and engage deeper layers of compiler logic, such as trait objects in Rust or templates in C++, both of which have received limited attention in prior studies (Yang et al., 2024b; Even-Mendoza et al., 2023; Ou et al., 2024; Wang and Jung, 2024). This limitation often results in missed opportunities to uncover subtle bugs that arise from intricate interactions between language features, which are prevalent in realistic programs (Wang et al., 2025).
@@
Our work.  We introduce Mut4All, a fully automated and language-agnostic framework for synthesizing mutation operators to support mutation-based compiler fuzzing. The core idea behind Mut4All is to automate the entire mutator lifecycle, from discovery and design to implementation, by coordinating specialized LLM agents that extract compiler-specific insights from historical bug reports in open-source repositories (e.g., GitHub). These real-world bug reports help identify bug-prone language features and provide concrete test programs that reveal compiler failures, allowing LLMs to focus on patterns that have previously caused actual bugs and produce more targeted and effective mutations. At a high level, Mut4All (1) analyzes real-world bug reports to identify language features prone to compiler bugs and formulates corresponding mutator specifications, (2) synthesizes concrete mutator implementations based on these specifications, and (3) validates and refines the generated mutators to ensure their correctness and effectiveness.
@@
  • Mutator Implementation Synthesis Agent: Synthesizes concrete mutator code guided by the specifications and high-quality AST modification examples. It is also fine-tuned with a curated set of high-quality seed mutators to produce more accurate and diverse transformations.
@@
  • Mutator Refinement Agent: Validates each synthesized mutator by applying it to bug-triggering programs collected from historical bug reports. If compilation or runtime errors occur, the agent automatically repairs it using error feedback. This process iterates until validation succeeds or a retry limit is reached.
@@
Following this workflow, we apply Mut4All to two prominent system-level programming languages, Rust and C++, and automatically analyze 1,000 bug reports (500 each for Rust and C++) from two open-source repositories: rustc for Rust and Clang for C++. In total, Mut4All synthesizes 319 mutators for Rust and 403 mutators for C++ without any human intervention, at an average cost of approximately \$0.08 per mutator using GPT-4o APIs. These operators target complex and error-prone language features, such as trait object layouts [^1] https://doc.rust-lang.org/book/ch10-02-traits.html in Rust and templates [^2] https://en.cppreference.com/w/cpp/language/templates.html in C++, which remain underexplored in prior work (Even-Mendoza et al., 2023; Ou et al., 2024; Sharma et al., 2023; Wang and Jung, 2024; Yang et al., 2024b).
@@
The synthesized mutators are then integrated into a customized fuzzer targeting production-level compilers, including rustc, gccrs, GCC, and Clang. Seeded with bug-triggering programs from historical reports and official test cases, the fuzzer uses crash and hang oracles, as well as differential testing across compilers of the same language (e.g., GCC and Clang for C++). We benchmark Mut4All against a range of state-of-the-art compiler fuzzers, including traditional fuzzers such as RustSmith (Sharma et al., 2023) and Rustlantis (Wang and Jung, 2024) for Rust, and CSmith (Yang et al., 2011), YARPGen (Livinskii et al., 2020; Li et al., 2023), and GrayC (Even-Mendoza et al., 2023) for C++. Additionally, we compare against recent LLM-based fuzzers, including Clozemaster (Gao et al., 2025) for Rust, as well as Fuzz4All (Xia et al., 2024) and MetaMut (Ou et al., 2024) for C++. The results indicate that Mut4All outperforms existing approaches by discovering more unique crashes, achieving the highest line coverage on Rust, and ranking second in coverage on C++. Particularly, compared with the approaches that directly leverage LLMs to mutate code (i.e., Clozemaster (Gao et al., 2025) and Fuzz4All (Xia et al., 2024)), our approach consistently performs better. In a 48-hour bug-hunting campaign, Mut4All uncovered a total of 96 bugs across both languages. For the Rust compilers (rustc and gccrs), Mut4All reported 62 bugs, of which 38 were confirmed as previously unknown and 7 had already been fixed by developers. For the C++ compilers (GCC and Clang), it reported 34 bugs, including 16 previously unknown and 1 already fixed. These results demonstrate that Mut4All can effectively synthesize high-quality mutators for complex language features, enabling practical bug discovery in real-world compilers.
@@
Contributions.  Mut4All addresses the key limitations of existing mutation-based compiler fuzzers. To overcome the limited scope and expressiveness of existing mutators (C1), it automatically invents new mutators guided by compiler-specific insights from open-source bug reports, often covering advanced language features. Unlike prior approaches that require manual effort to design mutators (C2) and are typically restricted to a single language or compiler (C3), Mut4All leverages LLMs as the generation engine—minimizing human intervention and enabling cross-language transfer through LLMs’ multilingual knowledge. In summary, this paper makes the following contributions:
@@
  • We propose Mut4All, the first fully automated multi-agent framework that synthesizes high-quality mutators for production-level compilers, guided by insights from historical compiler bug reports widely spread on open-source communities.
@@
  • We demonstrate the effectiveness of Mut4All by uncovering 96 compiler bugs (62 in rustc/gccrs, 34 in GCC/Clang), 54 of which have been confirmed as previously unknown, with 8 already fixed by developers. Additionally, Mut4All outperforms all baselines by finding more unique crashes and achieving the highest coverage on Rust and the second-highest on C++.
@@ 2.1. An Example for Rust
Figure 1. Rust mutator specification generated by Mut4All referring to Rust-124182.
We first demonstrate how Mut4All generates mutators for Rust by showcasing a confirmed bug it uncovered in the official Rust compiler (rustc). The process begins by collecting historical bug reports from the Rust compiler’s official GitHub repository, each typically including a descriptive title, a triggering code snippet, compiler output, and a stack trace, as shown in the top half of Figure 1. These reports provide valuable insights into the compiler’s behavior and highlight language features that are insufficiently supported and may lead to new bugs.
Given this bug report, Mut4All invokes a pipeline of three LLM agents to generate a new mutator. The first agent, Mutator Invention, analyzes the bug description and the triggering code to identify key language features responsible for the bug. It then synthesizes a mutator specification comprising a detailed description, syntax correctness constraints, and a concrete Before and After example illustrating the transformation. As shown in Figure 1, the historical bug (Rust-124182333https://github.com/rust-lang/rust/issues/124182) is recognized by Mut4All as related to generic parameter handling and type layout, which is based on the presence of a generic type in the code snippet and a reference to type layout in the bug title. This insight guides the LLM agent to invent a mutator that duplicates a generic type and wraps it inside a tuple, as detailed in the specification in the bottom half of Figure 1. Note that this mutator is much more sophisticated and creative than those generated by existing methods such as MetaMut (Ou et al., 2024), which typically rely on a fixed set of syntactic transformations. In contrast, our approach leverages insights extracted from real-world bug reports to design mutators that go well beyond the expressive capabilities of these existing techniques.
@@
Listing 1: The core implementation of the Rust mutator.
@@
Listing 2: The new bug Rust-14XX11 uncovered by Mut4All.
Next, the Mutator Implementation Synthesis agent, which is already enhanced via fine-tuning with a set of high-quality mutators, leverages the generated specification, our customized prompt templates, and several AST transformation examples extracted from Rust’s official AST library to produce a concrete mutator that performs the intended program transformation. The core logic of the resulting mutator is shown in Listing 1, which matches a generic type T at the AST level and replaces it with a tuple type (T, T). This transformation is designed to trigger compiler behavior related to generic types and type layout—the root cause of the original bug. In this case, the LLM generates a syntactically correct and semantically valid mutator without requiring any repair iterations. As a result, the Mutator Refinement agent is skipped for this example.
During our fuzzing phase, the mutator generates new program variants that stress the compiler’s type system in unexpected ways. As shown in Listing 2, Line 3 is mutated by introducing a self-referential tuple type VirtualWrapper<T>(T, T). This transformation causes the compiler to cast a pointer to the trait object dyn MyTrait into a pointer to the concrete type VirtualWrapper<T>, which is disallowed. The compiler fails to handle this edge case properly, resulting in an internal assertion failure. This bug, reported as Rust-14XX11, has been confirmed by Rust compiler developers as a previously unknown crash in rustc.
@@ 2.2. An Example for C++
Similarly, Mut4All generates mutators for C++ compilers by analyzing bug reports from LLVM’s GitHub repository. The process mirrors that of Rust, where LLM agents collaboratively invent, implement, and refine mutators based on compiler-specific insights. Figure 2 presents an example produced by Mut4All for C++.
The original bug report (LLVM-104802444https://github.com/llvm/llvm-project/issues/104802) describes a crash in the Clang C++ compiler triggered by a class that declares a member of its own type. In this case, the compiler fails to handle the invalid structure gracefully and enters infinite recursion during type checking. This suggests that recursive type patterns are not robustly guarded in certain compilation paths. Based on this insight, Mut4All’s Mutator Invention agent designs a mutator that intentionally injects self-referential type declarations, with the corresponding mutator specification shown in the bottom half of Figure 2. As mentioned above, existing methods such as MetaMut (Ou et al., 2024) are unable to design such mutators, as they rely on limited predefined syntactic transformations and do not account for language-specific features such as self-referential types.
Next, the Mutator Implementation Synthesis agent generates an implementation that inserts a self-referential field into a struct. Initially, the LLM produces a mutator using a deprecated AST API, resulting in a compiler error. After receiving the error feedback, Mut4All’s Mutator Refinement agent successfully repairs the mutator and produces a correct version (omitted here for brevity). The final mutator inserts a field such as S t; into a struct S, provoking recursive type behavior and potentially unsafe casting. Listing 3 shows an example of the mutated program, where the inserted self-reference interferes with a bit cast operation. This results in a compiler crash, reported as bug GCC-12XX98 which has been confirmed by GCC maintainers.
Figure 2. C++ mutator specification generated by Mut4All referring to LLVM-104802.

### v2

@@ Abstract.
In this work, we present Mut4All, a fully automated, cross-language framework that synthesizes mutation operators by leveraging LLM agents and compiler-specific insights extracted from bug reports on open-source compiler repositories. It is a multi-agent framework comprising three specialized LLM agents: (1) a mutator invention agent, which identifies mutation targets and generates mutator names, signatures, and descriptions based on compiler-related insights (e.g., language features explicitly mentioned in bug reports); (2) a mutator implementation synthesis agent, which is enhanced via fine-tuning to generate initial mutator implementations; and (3) a mutator refinement agent, which verifies and corrects the synthesized mutators using feedback from unit tests. To guide the LLM in using the correct and up-to-date AST modification APIs, we fine-tune the LLM of the latter two agents using a small set of hand-written mutators. Additionally, we propose an adaptive seed enhancement approach to generate more diverse seed programs. This pipeline analyzes 1000 bug reports (500 each for Rust and C++), and synthesizes 319 and 403 mutators for them. Each mutator only costs approximately \$0.08 using GPT-4o APIs. Using these mutators, our customized fuzzer has reported 96 bugs, and 58 of which have been confirmed as unknown bugs by the developers. There were 62 bugs in the Rust compilers (rustc and gccrs), 41 of which were confirmed as previously unknown, and 19 of which have already been fixed by developers. It has also reported 34 bugs in C++ compilers (GCC and Clang), including 17 confirmed as new, of which 3 have been fixed. Experimental results also show that Mut4All discovers a significant number of unique crashes, complementing existing fuzzers, and achieves the highest coverage on all Rust and C++ compilers.
Mutation-Based Compiler Fuzzing, LLM Agents, Compiler Bugs, Rust Compiler, C++ Compiler
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/XX/XXjournal: TOSEMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Compilers
@@ 1. Introduction
Compiler fuzzing (Chen et al., 2020; Marcozzi et al., 2019; Chen et al., 2016a) has proven effective in improving the reliability of production compilers such as GCC and rustc, uncovering numerous bugs over the past decades. Existing methods can be broadly categorized into two types: generation-based and mutation-based. Generation-based fuzzers, such as Csmith (Yang et al., 2011) and YARPGen (Livinskii et al., 2020, 2023), generate syntactically valid programs from scratch based on the compiler’s input grammar. However, ensuring the semantic validity of these test programs often requires substantial manual effort to design templates, particularly for modern programming languages with intricate features such as ownership and macros, which limits their scalability and adaptability.
In contrast, mutation-based fuzzers such as GrayC (Even-Mendoza et al., 2023), MetaMut (Ou et al., 2024), and TyMut (Wang et al., 2025b), which generate new test programs by applying mutation operators (i.e., mutators) to existing real-world programs (i.e., seed programs), are often preferred for compiler fuzzing, as they leverage existing code to avoid the need for designing complex templates and preserve both the syntactic and semantic validity of test programs. That being said, their effectiveness heavily depends on the quality and expressiveness of the mutators. Moreover, existing mutation-based fuzzers still face several major challenges in mutator design and implementation:
C1: Limited Expressiveness. Most existing mutation-based fuzzers define mutators as compositions of simple operators, such as arithmetic adjustments (e.g., replacing “+” with “-”) or control-flow rewrites (e.g., converting switch statements into if-else chains), as exemplified by GrayC (Even-Mendoza et al., 2023) and, to a more advanced extent, MetaMut (Ou et al., 2024). While effective at exposing bugs in basic language constructs, these mutators inherently lack expressiveness—the ability to construct program variants that incorporate advanced language features and engage in-depth of compiler logic, such as trait objects in Rust or templates in C++, both of which have received limited attention in prior studies (Yang et al., 2024c; Even-Mendoza et al., 2023; Ou et al., 2024; Wang and Jung, 2024; Liu et al., 2025). This limitation often results in missed opportunities to uncover subtle bugs that arise from intricate interactions between language features, which are prevalent in realistic programs (Wang et al., 2025a).
@@
Our work. We introduce Mut4All, a fully automated, cross-language framework for synthesizing mutation operators to support mutation-based compiler fuzzing. The core idea behind Mut4All is to automate the entire mutator lifecycle, from discovery and design to implementation, by coordinating multiple specialized LLM agents that extract compiler-specific insights from historical bug reports in open-source repositories (e.g., GitHub). These real-world bug reports help identify bug-prone language features and provide concrete test programs that reveal compiler failures, allowing LLMs to focus on patterns that have previously caused actual bugs and produce more targeted and effective mutations. At a high level, Mut4All (1) analyzes real-world bug reports to identify language features prone to compiler bugs and formulates corresponding mutator specifications, (2) synthesizes concrete mutator implementations based on these specifications, and (3) validates and refines the generated mutators to ensure their correctness and effectiveness. In addition, we propose an adaptive seed enhancement technique aimed at improving the diversity of the original seed corpus, which includes both historical bug reports and official test cases.
@@
  • Mutator Implementation Synthesis Agent: Synthesizes concrete mutator code guided by the specifications and high-quality few-shot mutator examples that illustrate AST modification codes. It is also fine-tuned with a curated set of high-quality seed mutators to produce more accurate and diverse transformations.
@@
  • Mutator Refinement Agent: Validates each synthesized mutator by applying it to bug-triggering programs collected from historical bug reports. If compilation or runtime errors occur, the agent automatically repairs them using error feedback. This process iterates until validation succeeds or a retry limit is reached.
@@
Following this workflow, we apply Mut4All to two prominent system-level programming languages, Rust and C++, and automatically analyze 1,000 bug reports (500 each for Rust and C++) from two open-source repositories: rustc for Rust and Clang for C++. To guide the LLM in using the correct and up-to-date AST manipulation APIs, we fine-tune the LLM of the last two agents using a small set of hand-written mutators as demonstrations. In total, Mut4All synthesizes 319 mutators for Rust and 403 mutators for C++ without any human intervention, at an average cost of approximately \$0.08 per mutator using GPT-4o APIs. These operators target complex and error-prone language features, such as trait object layouts in Rust [^1] https://doc.rust-lang.org/book/ch10-02-traits.html and templates in C++ [^2] https://en.cppreference.com/w/cpp/language/templates.html, which remain underexplored in prior work (Even-Mendoza et al., 2023; Ou et al., 2024; Wang et al., 2025b; Sharma et al., 2023; Wang and Jung, 2024; Yang et al., 2024c).
@@
The synthesized mutators are then integrated into a customized fuzzer targeting production-level compilers, including rustc, gccrs, GCC, and Clang. Seeded with the enhanced seed program set, Mut4All uses crash and hang oracles, as well as differential testing across compilers of the same language (e.g., GCC and Clang for C++). We benchmark Mut4All against a range of state-of-the-art compiler fuzzers, including traditional fuzzers such as RustSmith (Sharma et al., 2023) and Rustlantis (Wang and Jung, 2024) for Rust, and CSmith (Yang et al., 2011), YARPGen (Livinskii et al., 2020, 2023), GrayC (Even-Mendoza et al., 2023), and TyMut (Wang et al., 2025b) for C++. Additionally, we compare against the recent LLM-based fuzzers, including Clozemaster (Gao et al., 2025) for Rust, as well as Fuzz4All (Xia et al., 2024) and MetaMut (Ou et al., 2024) for C++. The results indicate that Mut4All outperforms existing approaches by discovering more unique crashes, achieving the highest line coverage on Rust and both C++ compilers. Particularly, compared with the approaches that directly leverage LLMs to mutate code (i.e., Clozemaster (Gao et al., 2025) and Fuzz4All (Xia et al., 2024)), our approach consistently performs better.
In the bug-hunting campaign, Mut4All uncovered a total of 96 bugs across both languages. For the Rust compilers (rustc and gccrs), Mut4All reported 62 bugs, of which 41 were confirmed as previously unknown and 19 had already been fixed by developers. For the C++ compilers (GCC and Clang), it reported 34 bugs, including 17 previously unknown ones and 3 of which had already been fixed. These results demonstrate that Mut4All can effectively synthesize high-quality mutators for complex language features, enabling practical bug discovery in real-world compilers.
@@
Contributions. Mut4All addresses the key limitations of existing mutation-based compiler fuzzers. To overcome the limited scope and expressiveness of existing mutators (C1), it automatically invents new mutators guided by compiler-specific insights from open-source bug reports, often covering advanced language features. Unlike prior approaches that require manual effort to design mutators (C2) and are typically restricted to a single language or compiler (C3), Mut4All leverages LLMs as the generation engine—minimizing human intervention and enabling cross-language transfer through LLMs’ multilingual knowledge. In summary, this work makes the following contributions:
@@
  • We propose Mut4All, the first fully automated multi-agent framework that synthesizes high-quality mutators for production-level compilers, guided by insights from historical compiler bug reports widely available on open-source communities.
  • We propose fine-tuning LLMs to guide the generation of correct AST modification APIs within mutators, and introduce an adaptive seed enhancement strategy to diversify the seed set, thereby improving the effectiveness of mutation-based compiler fuzzing.
@@
  • We demonstrate the effectiveness of Mut4All by uncovering 96 compiler bugs (62 in rustc/gccrs, 34 in GCC/Clang), 58 of which have been confirmed as previously unknown, with 22 already fixed by developers. Additionally, Mut4All outperforms all baselines by achieving the highest coverage on Rust and C++ compilers and recovering 13 unique crashes that all other fuzzers missed.
@@ 2.1. An Example for Rust
(a) Rust mutator specification via Rust-124182.
(b) C++ mutator specification via LLVM-104802.
Figure 1. Mutator specifications generated by Mut4All.
We first demonstrate how Mut4All generates mutators for Rust by showcasing a confirmed bug it uncovered in the official Rust compiler (rustc). The process begins by collecting historical bug reports from the Rust compiler’s official GitHub repository, each typically including a descriptive title, a triggering code snippet, compiler output, and a stack trace, as shown in the top half of Figure 1(a). These reports provide valuable insights into the compiler’s behavior and highlight language features that are insufficiently tested and may lead to new bugs.
Given this bug report, Mut4All invokes a pipeline of three LLM agents to generate a new mutator. The first agent, Mutator Invention, analyzes the bug description and the triggering code to identify key language features responsible for the bug. It then synthesizes a mutator specification comprising a detailed description, syntax well-formness constraints, and a concrete Before and After example illustrating the transformation. As shown in Figure 1(a), the historical bug (Rust-124182333https://github.com/rust-lang/rust/issues/124182) is recognized by Mut4All as related to generic parameter handling and type layout, which is based on the presence of a generic type in the code snippet and a reference to type layout in the bug title. This insight guides the LLM agent to invent a mutator that duplicates a generic type and wraps it inside a tuple, as detailed in the specification in the bottom half of Figure 1(a). Note that this mutator is much more sophisticated and creative than those generated by existing methods such as MetaMut (Ou et al., 2024), which typically rely on a fixed set of syntactic transformations. In contrast, our approach leverages insights extracted from real-world bug reports to design mutators that go well beyond the expressive capabilities of these existing techniques.
@@
Listing 1: The core implementation of the Rust mutator derived from Rust-124182.
@@
Listing 2: The new bug Rust-141911 uncovered by Mut4All.
Next, the Mutator Implementation Synthesis agent, leverages the generated specification, our customized prompt templates, and several AST transformation examples extracted from Rust’s official AST library to produce a concrete mutator that performs the intended program transformation. The backend LLM of this agent is already enhanced via fine-tuning with a set of high-quality human-written mutators across all target languages. The core logic of the resulting mutator is shown in Listing 1, which matches a generic type T at the AST level and replaces it with a tuple type (T, T). This transformation is designed to trigger compiler behavior related to generic types and type layout—the root cause of the original bug. In this case, the LLM generates a syntactically correct and semantically valid mutator without requiring any repair iterations. As a result, the Mutator Refinement agent is skipped for this example.
During our fuzzing phase, the mutator generates new program variants that stress-test the compiler’s type system in unexpected ways. As shown in Listing 2, Line 2 is mutated by introducing a self-referential tuple type VirtualWrapper<T>(T, T). This transformation causes the compiler to cast a pointer to the trait object dyn MyTrait into a pointer to the concrete type VirtualWrapper<T>, which is disallowed. The compiler fails to handle this edge case properly, resulting in an internal assertion failure. This bug, reported as Rust-141911, has been confirmed by Rust compiler developers as a previously unknown crash in rustc.
@@ 2.2. An Example for C++
Similarly, Mut4All generates mutators for C++ compilers by analyzing bug reports from LLVM’s GitHub repository. The process mirrors that of Rust, where LLM agents collaboratively invent, implement, and refine mutators based on compiler-specific insights. Figure 1(b) presents an example produced by Mut4All for C++.
The original bug report (LLVM-104802444https://github.com/llvm/llvm-project/issues/104802) describes a crash in the Clang C++ compiler triggered by a class that declares a member of its own type. In this case, the compiler fails to handle the invalid structure gracefully and enters infinite recursion during type checking. This suggests that recursive type patterns are not robustly guarded in certain compilation paths. Based on this insight, Mut4All’s Mutator Invention agent designs a mutator that intentionally injects self-referential type declarations, with the corresponding mutator specification shown in the bottom half of Figure 1(b). As mentioned above, existing methods such as MetaMut (Ou et al., 2024) are unable to design such mutators, as they rely on limited predefined syntactic transformations and do not account for language-specific features such as self-referential types.
Next, the Mutator Implementation Synthesis agent generates an implementation that inserts a self-referential field into a struct. Initially, the LLM produces a mutator using a deprecated AST API, resulting in a compiler error. After receiving the error feedback, Mut4All’s Mutator Refinement agent successfully repairs the mutator and produces a correct version, shown as Listing 3. The final mutator introduces recursive types, such as introducing a field of type S (i.e., S t;) into the struct, provoking recursive type behavior and potentially unsafe casting. Listing 4 shows an example of the mutated program, where the inserted self-reference interferes with a bit cast operation. This results in a compiler crash, reported as bug GCC-120798, which has been confirmed by GCC maintainers.
1void MutatorFrontendAction_6::Callback::run(const MatchFinder::MatchResult &Result) {
2 //Check whether the matched AST node is the target node
3 if (auto *CD = Result.Nodes.getNodeAs<clang::CXXRecordDecl>("ClassDecl")) {
4 //Filter nodes in header files
5 if (!CD $|$ !Result.Context->getSourceManager().isWrittenInMainFile(CD->getLocation()))
6 return;
7 if (!CD->isCompleteDefinition())
8 return;
9 //Get the source code text of target node
10 auto className = CD->getNameAsString();
11 if (className.empty())
12 return;
13 //Perform mutation on the source code text by applying string replacement
14 auto classText = stringutils::rangetoStr(*(Result.SourceManager), CD->getSourceRange());
15 std::string memberDecl = className + " self_member;";
[…]
17 classText = classText.insert(classText.rfind("}")-1, memberDecl);
18 //Replace the original AST node with the mutated one
19 Rewrite.ReplaceText(CharSourceRange::getTokenRange(CD->getSourceRange()), classText);
20 }
21}
Listing 3: The core implementation of the c++ mutator derived from LLVM-104802.